OpenStudy (anonymous):

At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party?

OpenStudy (anonymous):

12 people

OpenStudy (anonymous):

ok thats easy do you know combinations?

OpenStudy (anonymous):

irst person shakes hand with everyone else: n-1 times(n-1 persons) second person shakes hand with everyone else(not with 1st as its already done): n-2 times 3rd person shakes hands with remaining persons: n-3 So total handshakes will be = (n-1) + (n-2) + (n-3) +…… 0; = (n-1)*(n-1+1)/2 = (n-1)*n/2 = 66 = n^2 -n = 132 =(n-12)(n+11) = 0; = n = 12 OR n =-11 -11 is ruled out so the answer is 12 persons.

OpenStudy (anonymous):

Hmm interesting way of solving the problem. I did nC2=66 and got the same thing. That is n(n-1)=132 solve for n

OpenStudy (anonymous):

i did \(1+2+3+4+5+6+7+8+9+10+11\) and got the same answer more than one way to skin a cat

OpenStudy (anonymous):

good job lol!